OpenStudy (anonymous):

What are the solutions of the compound inequality? Graph the solutions. –2 < 4x – 10 < 6

OpenStudy (callisto):

–2 < 4x – 10 < 6 1. add 10 to all the 'terms' like this –2 + 10 < 4x – 10 +10 < 6 + 10 What do you get?

OpenStudy (anonymous):

8<4x0<16

OpenStudy (callisto):

8<4x<16 ... Now, divide all the 'terms' by 4, like this 8/4 <4x/4 <16/4 What do you get?

OpenStudy (anonymous):

2<x<4

OpenStudy (callisto):

That's the answer, I suppose :P Now, do you know what/how should you graph it? number line? or ..?

OpenStudy (anonymous):

yes i do:) | x | + 10 = 1 will u help me with this?:)

OpenStudy (callisto):

Probably the last question I can help... first, subtract 10 from both sides. what do you get?

OpenStudy (anonymous):

i will only have one more after this..... and x+0=-9

OpenStudy (callisto):

You need to know that x+0 =x , so you don't have to write '+0' or '-0' and also you need to keep the absolute sign... |x| = -9 Absolute value means that we just consider the +ve form of a number. For example x = 3 |x| = |3| = 3 x=-3 |x| = |-3| = 3 In this case |x| = -9, which does not make sense...

OpenStudy (anonymous):

so whats the answer

OpenStudy (callisto):

No solution :| Or you can check the question again to see if |x| + 10 =1 is really your question..

OpenStudy (anonymous):

ok this is my last one What are the solutions of ? Write the solutions as either the union or the intersection of two sets.

OpenStudy (anonymous):

What are the solutions of ? Write the solutions as either the union or the intersection of two sets. A. B. C. D.

OpenStudy (callisto):

Hmm... you missed something in the question...

OpenStudy (anonymous):

x+6 is greater then or = to 5 sorry;)

OpenStudy (callisto):

Never mind :) Let this sign ≧ be greater than or equal to ... x+6 ≧ 5 Subtract 6 from both sides, what do you get?

OpenStudy (anonymous):

x(ge)-1

OpenStudy (anonymous):

\[\ge\]

OpenStudy (callisto):

Yes. Finished :D

OpenStudy (anonymous):

thank you

OpenStudy (callisto):

welcome~